A parachutist's rate during a free fall reaches 165

 feet per second. What is this rate in meters per second? At this rate, how many meters will the parachutist fall during 15 seconds of free fall? In your computations, assume that 1 meter is equal to 3.3 feet. Do not round your answers.


1
Expert's answer
2021-01-25T00:12:24-0500

Assuming that:


"1 m=3.3 ft"


We can make the conversion from feet per second to meters per second with this procedure:


"(165\ \frac{ft}{s})(\frac{1\ m}{3.3\ ft})=50\ \frac{m}{s}"

Let "x" be the amount of meters the parachutist will fall during 15 seconds of free fall


Having that in 1 second the parachutist falls 50 meters, we can calculate "x" :


"x=(50\\ \\frac{m}{s})(15\\ s)\n\u200b"



x=750 m